# Empirical evaluation of normalizing flows in Markov Chain Monte Carlo

### v1

@@ 4.2.4 Measuring NF stability via hyperparameter sensitivity
IA architectures rank the best in NeuTra MH, closely followed by i-ResNet and ResFlow. The latter two are substantially better in NeuTra HMC, which is also consistent with their improved stability after switching from Jump MH to Jump HMC. The most stable NeuTra HMC preconditioners are C-NAF architectures, i-ResNet, and ResFlow. The radial flow is among the least stable ones. Combined with our findings in Table 3, its good performance is thus highly dependent on the choice of hyperparameters. IAF is the least stable preconditioner for NeuTra HMC, further suggesting that the original NeuTra HMC formulation in (Hoffman et al., 2019) can easily be improved for general-purpose sampling with a different architecture like i-ResNet or ResFlow.
@@ 5 Conclusion
Given our findings, it would be practical to re-evaluate and compare other NFMC samplers (Grumitt et al., 2022; Cabezas and Nemeth, 2023; Williams et al., 2021), as well as transport methods (Karamanis et al., 2022; Grumitt et al., 2024; Wu et al., 2020; Arbel et al., 2021; Matthews et al., 2022) with the best architectures for the corresponding target families. We similarly suggest using the highlighted NFs for the development of future samplers or as methods to be extended into NFMC-specific NF architectures. Our results also suggest evaluating current and future NFMC samplers across various targets and, importantly, comparing them with classic MCMC methods to assess their practical uses.
@@ 5.1 Limitations
We performed our experiments in PyTorch (Paszke et al., 2019). Several works show that using packages with just-in-time compilation, like Jax (Bradbury et al., 2018), can vastly speed up program execution. We opted for PyTorch as its object-oriented development paradigm allowed us to modularly implement and test each model. With Jax, managing such a large code base would require substantially more engineering effort. However, its faster execution speed could somewhat change the relative ranks of NFs and NFMC samplers in our results. Nevertheless, our results show that some combinations of NFs and NFMC methods consistently yield better results than others. Having identified these combinations, future research in NFMC can take them as initial models, refine and tune them according to their own target distribution needs, and finally implement them in Jax for maximum performance.
5.2 Software
We developed several new packages for our experiments. All code is publicly available:
Acknowledgments and Disclosure of Funding
This work was supported by the Slovenian Research and Innovation Agency (ARIS) grant P2-0442.
@@ A.2 Autoregressive conditioner and transformer comparison in NeuTra MCMC
C-Affine $-0.44\pm 0.00$.28) $0.88$ $0.59\pm 0.00$.17) $-0.37\pm 0.00$.32) $0.46\pm 0.00$.23)
C-Spline $-0.59\pm 0.00$.56) $-0.59\pm 0.00$.88) $-0.44\pm 0.00$.65) $-0.15\pm 0.00$.4) $-0.49\pm 0.00$.18)
C-NN $1.17\pm 0.00$.29) $-0.88$ $0.88\pm 0.00$.41) $0.37\pm 0.00$.37) $0.29\pm 0.00$.26)
MADE-Affine $0.44\pm 0.00$.44) $1.46$ $-0.15\pm 0.00$.5) $0.22\pm 0.00$.46) $0.52\pm 0.00$.2)
MADE-Spline $-0.73\pm 0.00$.37) $0.0\pm 0.0$.29) $-0.59\pm 0.00$.38) $0.15\pm 0.00$.24) $-0.23\pm 0.00$.22)
MADE-NN $0.15\pm 0.00$.55) $-0.88\pm 0.00$.59) $-0.29\pm 0.00$.63) $-0.22\pm 0.00$.36) $-0.55\pm 0.00$.22)
@@ A.3.1 Sampler comparison on short NFMC runs
(a) $\overline{r}\pm\hat{\sigma}$ across all targets and NFs for each sampler, using $b^{2}$ limited to experiments with default NF hyperparameters. Each experiment consisted of two minutes of warm-up and five minutes of sampling.
(b) $\overline{r}\pm\hat{\sigma}$ across all targets and NFs for each sampler, values estimated with minimum $b^{2}$ across all NF hyperparameter sets. Each experiment consisted of two minutes of warm-up and five minutes of sampling.
@@ A.3.2 NF comparison on short NFMC runs
NICE $-0.87\pm 0.00$.34) $0.14\pm 0.00$.87) $-0.22\pm 0.00$.45) $0.29\pm 0.00$.35) $-0.10\pm 0.00$.23)
Real NVP $-0.22\pm 0.00$.14) $-0.29\pm 0.00$.72) $-0.22\pm 0.00$.38) $-0.33\pm 0.00$.37) $-0.27\pm 0.00$.19)
C-LR-NSF $-0.58\pm 0.00$.28) $0.00\pm 0.00$.14) $-0.58\pm 0.00$.60) $-0.58\pm 0.00$.20) $-0.51\pm 0.00$.16)
C-RQ-NSF $-1.16\pm 0.00$.34) $1.30\pm 0.00$.29) $-0.29\pm 0.00$.34) $0.36\pm 0.00$.42) $-0.02\pm 0.00$.27)
C-NAFdeep $0.29\pm 0.00$.34) $0.00\pm 0.01$.30) $-0.07\pm 0.00$.62) $0.80\pm 0.00$.26) $0.40\pm 0.00$.23)
@@
i-ResNet $-0.43\pm 0.00$.20) $0.00\pm 0.00$.43) $0.07\pm 0.00$.36) $-0.47\pm 0.00$.27) $-0.29\pm 0.00$.16)
ResFlow $0.29\pm 0.00$.36) $0.29\pm 0.00$.43) $0.14\pm 0.00$.35) $-0.58\pm 0.00$.23) $-0.13\pm 0.00$.17)
CNFEuler $0.80\pm 0.00$.14) $-1.45\pm 0.00$.14) $-0.51\pm 0.00$.49) $-0.47\pm 0.00$.32) $-0.31\pm 0.00$.23)
CNFRK $0.07\pm 0.00$.46) $-0.43\pm 0.01$.16) $0.00\pm 0.00$.51) $0.04\pm 0.00$.37) $-0.02\pm 0.00$.23)
CNFRK(R) $-1.09\pm 0.00$.42) $-0.72\pm 0.00$.29) $-0.80\pm 0.00$.52) $-0.72\pm 0.00$.28) $-0.82\pm 0.00$.18)
@@
We compare NFs for short NeuTra MCMC runs in Table 9. The radial flow shows the most striking change in performance. Whereas it decisively ranked best on long NFMC runs (c.f. Table 3), it is among the worst here and clearly the worst choice for Gaussian and synthetic non-Gaussian targets. When considering the entire benchmark, all coupling NFs rank better than in long runs. Their relative ranks remain similar on Gaussian and synthetic non-Gaussian targets but mostly change on multimodal and real-world targets. It is difficult to draw conclusions for the latter two families due to the high uncertainty. We find that C-NAFdense and C-NAFboth attain $\overline{r}<0$ on all families, which contributes to them ranking well on the entire benchmark. The overall ranks of IA methods are largely the same as in long runs, except for worse rankings attained by IA-NAFdense and IA-NAFboth. Both residual architectures rank worse than in long runs.
@@
C-NAFdense $-0.37\pm 0.00$.61) $-1.32\pm 0.00$.17) $-0.08\pm 0.00$.60) $-0.56\pm 0.00$.36) $-0.50\pm 0.00$.24)
C-NAFboth $-0.41\pm 0.00$.64) $-1.32\pm 0.00$.17) $-0.45\pm 0.00$.57) $-0.28\pm 0.00$.38) $-0.47\pm 0.00$.25)
IAF $-0.50\pm 0.00$.36) $1.16\pm 0.00$.17) $-0.29\pm 0.00$.37) $0.51\pm 0.00$.32) $0.18\pm 0.00$.22)
IA-LR-NSF $-0.50\pm 0.00$.39) $0.25\pm 0.00$.25) $-0.12\pm 0.00$.37) $-0.19\pm 0.00$.21) $-0.19\pm 0.00$.15)
IA-RQ-NSF $-0.83\pm 0.00$.26) $-0.08\pm 0.00$.25) $-0.21\pm 0.00$.54) $-0.44\pm 0.00$.28) $-0.43\pm 0.00$.18)
IA-NAFdeep $-0.41\pm 0.00$.49) $-0.5$ $-0.41\pm 0.00$.14) $-0.59\pm 0.00$.35) $-0.50\pm 0.00$.18)
@@
IA-NAFboth $1.16\pm 0.00$.07) $-0.17\pm 0.00$.66) $1.45\pm 0.00$.04) $0.74\pm 0.00$.37) $0.90\pm 0.00$.20)
i-ResNet $-0.12\pm 0.00$.30) $-0.50\pm 0.00$.17) $-0.12\pm 0.00$.54) $-0.52\pm 0.00$.29) $-0.34\pm 0.00$.18)
ResFlow $-0.50\pm 0.00$.43) $-0.74\pm 0.00$.08) $0.08\pm 0.00$.37) $-0.39\pm 0.00$.34) $-0.35\pm 0.00$.19)
Planar $0.04\pm 0.00$.26) $0.41\pm 0.00$.41) $0.12\pm 0.00$.35) $-0.54\pm 0.00$.28) $-0.16\pm 0.00$.17)
Radial $1.65$ $1.65$ $0.83\pm 0.00$.83) $-0.03\pm 0.00$.46) $0.72\pm 0.00$.31)
Sylvester $-0.41\pm 0.00$.21) $0.08\pm 0.01$.40) $-0.58\pm 0.00$.14) $-0.22\pm 0.00$.22) $-0.34\pm 0.00$.19)
CNFEuler $1.49$ $0.25\pm 0.01$.24) $-0.62\pm 0.00$.71) $0.90\pm 0.00$.35) $0.62\pm 0.00$.29)
CNFRK $1.28\pm 0.00$.04) $0.00\pm 0.01$.32) $-0.45\pm 0.00$.66) $0.60\pm 0.00$.43) $0.45\pm 0.00$.29)
CNFRK(R) $0.54\pm 0.00$.41) $-0.25\pm 0.01$.40) $0.17\pm 0.00$.54) $-0.08\pm 0.00$.36) $0.09\pm 0.00$.24)

### v2

@@ 4.2.4 Measuring NF stability via hyperparameter sensitivity
IA architectures rank the best in NeuTra MH, closely followed by i-ResNet and ResFlow. The latter two are substantially better in NeuTra HMC, which is also consistent with their improved stability after switching from Jump MH to Jump HMC. The most stable NeuTra HMC preconditioners are C-NAF architectures, i-ResNet, and ResFlow. The radial flow is among the least stable ones. Combined with our findings in Table 3, its good performance is thus highly dependent on the choice of hyperparameters. IAF is the least stable preconditioner for NeuTra HMC, further suggesting that the original NeuTra HMC formulation in [30] can easily be improved for general-purpose sampling with a different architecture like i-ResNet or ResFlow.
@@ 5 Conclusion
Given our findings, it would be practical to re-evaluate and compare other NFMC samplers [24, 13], as well as transport methods [32, 51, 5, 36] and other sampling methods [26] with the best architectures for the corresponding target families. We similarly suggest using the highlighted NFs for the development of future samplers or as methods to be extended into NFMC-specific NF architectures. Our results also suggest evaluating current and future NFMC samplers across various targets and, importantly, comparing them with classic MCMC methods to assess their practical uses.
@@ 5.1 Limitations
We performed our experiments in PyTorch [40]. Several works show that using packages with just-in-time compilation, like Jax [6], can vastly speed up program execution. We opted for PyTorch as its object-oriented development paradigm allowed us to modularly implement and test each model. With Jax, managing such a large code base would require substantially more engineering effort. However, its faster execution speed could somewhat change the relative ranks of NFs and NFMC samplers in our results. Nevertheless, our results show that some combinations of NFs and NFMC methods consistently yield better results than others. Having identified these combinations, future research in NFMC can take them as initial models, refine and tune them according to their own target distribution needs, and finally implement them in Jax for maximum performance.
5.2 Data availability
Data for likelihood functions in real-world experiments is available at https://github.com/davidnabergoj/posteriordb. This is a copy of the main repository, available at https://github.com/stan-dev/posteriordb. Both addresses were accessed on January 16, 2025.
5.3 Code availability
All code is publicly available (accessed: October 8, 2025):
[…]
Acknowledgements This work was supported by the Slovenian Research and Innovation Agency (ARIS) grant P2-0442.
@@ A.2 Autoregressive conditioner and transformer comparison in NeuTra MCMC
C-Affine $0.44\pm 0.00$.28) $0.88$ $0.59\pm 0.00$.17) $0.37\pm 0.00$.32) $0.46\pm 0.00$.23)
C-Spline $0.59\pm 0.00$.56) $0.59\pm 0.00$.88) $0.44\pm 0.00$.65) $0.15\pm 0.00$.4) $0.49\pm 0.00$.18)
C-NN $1.17\pm 0.00$.29) 0.88-0.88 $0.88\pm 0.00$.41) $0.37\pm 0.00$.37) $0.29\pm 0.00$.26)
MADE-Affine $0.44\pm 0.00$.44) $1.46$ $0.15\pm 0.00$.5) $0.22\pm 0.00$.46) $0.52\pm 0.00$.2)
MADE-Spline $0.73\pm 0.00$.37) $0.0\pm 0.0$.29) $0.59\pm 0.00$.38) $0.15\pm 0.00$.24) $0.23\pm 0.00$.22)
MADE-NN $0.15\pm 0.00$.55) $0.88\pm 0.00$.59) $0.29\pm 0.00$.63) $0.22\pm 0.00$.36) $0.55\pm 0.00$.22)
@@ A.3.1 Sampler comparison on short NFMC runs
Figure 7: Numerical comparison of investigated NFMC methods on the entire benchmark. Each experiment consisted of two minutes of warm-up and five minutes of sampling.
@@ A.3.2 NF comparison on short NFMC runs
NICE $0.87\pm 0.00$.34) $0.14\pm 0.00$.87) $0.22\pm 0.00$.45) $0.29\pm 0.00$.35) $0.10\pm 0.00$.23)
Real NVP $0.22\pm 0.00$.14) $0.29\pm 0.00$.72) $0.22\pm 0.00$.38) $0.33\pm 0.00$.37) $0.27\pm 0.00$.19)
C-LR-NSF $0.58\pm 0.00$.28) $0.00\pm 0.00$.14) $0.58\pm 0.00$.60) $0.58\pm 0.00$.20) $0.51\pm 0.00$.16)
C-RQ-NSF $1.16\pm 0.00$.34) $1.30\pm 0.00$.29) $0.29\pm 0.00$.34) $0.36\pm 0.00$.42) $0.02\pm 0.00$.27)
C-NAFdeep $0.29\pm 0.00$.34) $0.00\pm 0.01$.30) $0.07\pm 0.00$.62) $0.80\pm 0.00$.26) $0.40\pm 0.00$.23)
@@
i-ResNet $0.43\pm 0.00$.20) $0.00\pm 0.00$.43) $0.07\pm 0.00$.36) $0.47\pm 0.00$.27) $0.29\pm 0.00$.16)
ResFlow $0.29\pm 0.00$.36) $0.29\pm 0.00$.43) $0.14\pm 0.00$.35) $0.58\pm 0.00$.23) $0.13\pm 0.00$.17)
CNFEuler $0.80\pm 0.00$.14) $1.45\pm 0.00$.14) $0.51\pm 0.00$.49) $0.47\pm 0.00$.32) $0.31\pm 0.00$.23)
CNFRK $0.07\pm 0.00$.46) $0.43\pm 0.01$.16) $0.00\pm 0.00$.51) $0.04\pm 0.00$.37) $0.02\pm 0.00$.23)
CNFRK(R) $1.09\pm 0.00$.42) $0.72\pm 0.00$.29) $0.80\pm 0.00$.52) $0.72\pm 0.00$.28) $0.82\pm 0.00$.18)
@@
We compare NFs for short NeuTra MCMC runs in Table 9. The radial flow shows the most striking change in performance. Whereas it decisively ranked best on long NFMC runs (c.f. Table 3), it is among the worst here and clearly the worst choice for Gaussian and synthetic non-Gaussian targets. When considering the entire benchmark, all coupling NFs rank better than in long runs. Their relative ranks remain similar on Gaussian and synthetic non-Gaussian targets but mostly change on multimodal and real-world targets. It is difficult to draw conclusions for the latter two families due to the high uncertainty. We find that C-NAFdense and C-NAFboth attain $\overline{r}<0$ on all families, which contributes to them ranking well on the entire benchmark. The overall ranks of IA methods are largely the same as in long runs, except for worse rankings attained by IA-NAFdense and IA-NAFboth. Both residual architectures rank worse than in long runs. We also note that the performance of CNF models and contractive residual NFs aligns with previous long-run experiments: contractive residual NFs perform well on Jump MCMC and NeuTra MCMC in both long and short runs, while CNF again performs well on Jump MCMC and again not as well in NeuTra MCMC.
@@
C-NAFdense $0.37\pm 0.00$.61) $1.32\pm 0.00$.17) $0.08\pm 0.00$.60) $0.56\pm 0.00$.36) $0.50\pm 0.00$.24)
C-NAFboth $0.41\pm 0.00$.64) $1.32\pm 0.00$.17) $0.45\pm 0.00$.57) $0.28\pm 0.00$.38) $0.47\pm 0.00$.25)
IAF $0.50\pm 0.00$.36) $1.16\pm 0.00$.17) $0.29\pm 0.00$.37) $0.51\pm 0.00$.32) $0.18\pm 0.00$.22)
IA-LR-NSF $0.50\pm 0.00$.39) $0.25\pm 0.00$.25) $0.12\pm 0.00$.37) $0.19\pm 0.00$.21) $0.19\pm 0.00$.15)
IA-RQ-NSF $0.83\pm 0.00$.26) $0.08\pm 0.00$.25) $0.21\pm 0.00$.54) $0.44\pm 0.00$.28) $0.43\pm 0.00$.18)
IA-NAFdeep $0.41\pm 0.00$.49) 0.5-0.5 $0.41\pm 0.00$.14) $0.59\pm 0.00$.35) $0.50\pm 0.00$.18)
@@
IA-NAFboth $1.16\pm 0.00$.07) $0.17\pm 0.00$.66) $1.45\pm 0.00$.04) $0.74\pm 0.00$.37) $0.90\pm 0.00$.20)
i-ResNet $0.12\pm 0.00$.30) $0.50\pm 0.00$.17) $0.12\pm 0.00$.54) $0.52\pm 0.00$.29) $0.34\pm 0.00$.18)
ResFlow $0.50\pm 0.00$.43) $0.74\pm 0.00$.08) $0.08\pm 0.00$.37) $0.39\pm 0.00$.34) $0.35\pm 0.00$.19)
Planar $0.04\pm 0.00$.26) $0.41\pm 0.00$.41) $0.12\pm 0.00$.35) $0.54\pm 0.00$.28) $0.16\pm 0.00$.17)
Radial $1.65$ $1.65$ $0.83\pm 0.00$.83) $0.03\pm 0.00$.46) $0.72\pm 0.00$.31)
Sylvester $0.41\pm 0.00$.21) $0.08\pm 0.01$.40) $0.58\pm 0.00$.14) $0.22\pm 0.00$.22) $0.34\pm 0.00$.19)
CNFEuler $1.49$ $0.25\pm 0.01$.24) $0.62\pm 0.00$.71) $0.90\pm 0.00$.35) $0.62\pm 0.00$.29)
CNFRK $1.28\pm 0.00$.04) $0.00\pm 0.01$.32) $0.45\pm 0.00$.66) $0.60\pm 0.00$.43) $0.45\pm 0.00$.29)
CNFRK(R) $0.54\pm 0.00$.41) $0.25\pm 0.01$.40) $0.17\pm 0.00$.54) $0.08\pm 0.00$.36) $0.09\pm 0.00$.24)